### v1

@@ Abstract
We present ZipNN a lossless compression tailored to neural networks. Somewhat surprisingly, we show that specific lossless compression can gain significant network and storage reduction on popular models, often saving $33\%$ and at times reducing over $50\%$ of the model size. We investigate the source of model compressibility and introduce specialized compression variants tailored for models that further increase the effectiveness of compression. On popular models (e.g. Llama 3) ZipNN shows space savings that are over $17\%$ better than vanilla compression while also improving compression and decompression speeds by $62\%$. We estimate that these methods could save over an ExaByte per month of network traffic downloaded from a large model hub like Hugging Face.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
With scale, we have learned that models gain stronger abilities and with it, gain popularity. With scale, models also require more storage space and memory and with popularity more communication bandwidth. Taken together, we observe strains on communication bottlenecks that call for efficient solutions. Storage requirements, while often ignored, may accumulate to hundreds or thousands of times the size of a model if checkpoints Biderman et al. (2023) or distributed updates are to be saved (c.f., 6) Kandpal et al. (2023); Don-Yehiya et al. (2023); Zhang et al. (2021).
Similarly, models are repeatedly moved around in multiple channels: from a storage hub to inference machines; from training/fine-tuning nodes to the storage backend; between GPU nodes during distributed training, and so on. Network hubs epitomize the strains by model size. For instance, with over 14.5 GBs and 2.77 M downloads per month from Hugging Face Wolf et al. (2019) Mistral Jiang et al. (2023) alone requires 40 PBs of transferred information a month.
A large body of work has been aimed at reducing model sizes focusing on the parameter sizes and the number of computations in inference. Such methods transform the model into a smaller one in an irreversible fashion. For example, distillation Gou et al. (2021), pruning Ma et al. (2021) and quantization Gholami et al. (2021) all lose some information and, potentially, performance, for improved efficiency. Since these methods’ main focus is on inference speed, they are bound to the format of an actual running model. As such, they don’t necessarily push the space-saving to its limits, and are not stored in the minimal possible way.
@@ 1 Introduction
Surprisingly, we observe (§3) that even standard lossless compressors like zlib Deutsch & Gailly (1996) or zstd Collet & Kucherawy (2018) can achieve non-negligible savings. While common rationale expects model parameters to have high entropy and therefore be non-compressible, we find that in reality there is less entropy than what the representation offers. Our goal is therefore to find a compression method that maximizes compression benefits for models while also being aware of compression and decompression speeds.
We identify the source of model compressibility as the floating point range that actually exists in models. Specifically, we find that the exponent component in a floating point parameter is highly skewed and therefore very compressible. To this end, we devise a compression method that separates the exponent bits from the rest of the bits that often show no compressibility. We also identified that the source of compressibility of the exponent is entirely the skewed distribution of single bytes. This means that compressors that search for multi-byte repetitions are both unnecessarily time-consuming and unhelpful in this case, and in particular all of the Lempel-Ziv algorithms Ziv & Lempel (1977; 1978) hardly achieve any data reduction. Instead, we use only Entropy Encoding and specifically Huffman codes Huffman (1952), improving both performance and compression ratio in doing so. Another speed improvement comes from identifying the non-compressible parts and avoiding time-consuming compression attempts on them. The observations described above are relevant to most of the models, but not all of them. ZipNN therefore identifies the characteristics of the model at hand and determines which strategy best suits it.
@@
We classify popular models into categories with distinct compressibility traits. We highlight the clean models category, where models have undergone some rounding after the training phase. The rounding allows compression of the other bits (apart for the exponent) and therefore clean models achieve effective compression, at times reaching 55% savings. For example, the popular RoBERTa model Liu et al. (2019b) falls into this category. Our compressor is therefore challenged with the task of inline identifying the potential space savings in such a model and compressing it accordingly.
When comparing ZipNN to a state-of-the art compressor like Zstd we get the following improvements: On BF16 models our compressor achieves a 17% improvement in compression ratio and a 62% speedup in compression/decompression. For clean models, this is even larger with a 34% improvement in space and a 4.6X speedup in compression time and 83% speedup in decompression.
@@
2 Background
2.1 Motivation - use cases
With small models weighing about a Gigabyte Devlin et al. (2019) and large ones Terrabytes Fedus et al. (2021) storage and network overheads become an issue for many purposes. Moreover, common use cases require many model types or model versions and hence increased resources. We list some below as a motivation.
Table 1: Top ranked downloaded models from Hugging Face.
@@ 2.1.1 Model Hubs
Large model repositories or hubs like Hugging Face Wolf et al. (2019), Model Zoo Yu Koh (2018), PyTorch Pytorch (2019), Tensorflow Google (2018), Adapter Pfeiffer et al. (2020), IBM watsonx.data watsonx.data and Qualcomm® AI Qualcomm AI Hub hold a large number of models and serve numerous download requests of popular models. Hugging Face, the largest of these hubs, revealed in a statement in August 2024, that it holds 1.3M models, with a cumulative storage space of 12PB. They also serve 1 billion daily requests amounting to a network bandwidth of around 6 PetaBytes per day! Table 1 shows some of the top-rankedmodels [^1] Rank as of October 2024. and their compression ratio (using the methods described in Section 2.3). As seen, the potential traffic savings from compression is substantial.
@@
In this use case, there are three ways in which compression can be beneficial, the first and the most important is to reduce the amount of data transferred, the second is to reduce the amount of data stored and the third is to reduce the time to download and upload those models.
2.1.2 Distributed and Decentralized Training
@@
During training of large models, training nodes transfer data between them to overcome the need to save the full model and computation on a single GPU/node. In some methods, only the model weights are transferred between nodes, and in other methods, the optimizer weights and gradients are transferred as well Zhao et al. (2023). Either way, distributed training is usually limited by data transfer between nodes, an issue compression addresses directly. For example: FSDP (Fully Sharded Data Parallel) is a PyTorch method that during training, in addition to transferring model weights, also transfers gradients and optimizer states, causing a networking bottleneck that limits the training of large models.
@@
Another training paradigm is decentralized or federated training, which proposes accommodating training of the same model by different contributors. This ranges from federated learning that contributes gradients Zhang et al. (2021), to contributing partially trained models Li et al. (2022), from changing the kinds of updates done Lialin et al. (2023) to relying on volunteer computing Diskin et al. (2021), or even relying on different objectives and expertise all merged into the same model Don-Yehiya et al. (2023). All of these methods inherently transfer and store numerous model versions, which also drove dedicated version control frameworks Kandpal et al. (2023).
2.1.3 Checkpoints and Versions
@@
During model creation, multiple intermediate versions of the models are commonly saved. This often includes tests on the training regime such as hyperparameter tuning Turner et al. (2021). Even during the training of a single model, the current model is periodically checkpointed to recover after a crash, to select the best checkpoint from a few options Dodge et al. (2020), for analysis Biderman et al. (2023), improve performance Junczys-Dowmunt et al. (2018); Sandler et al. (2023), etc. Even though saving during checkpointing rarely slows the training time, it does burden the networking and storage, limiting the frequency and amount of saved and shared checkpoints, which are encouraged by the community (e.g.; Biderman et al. (2023); Liu et al. (2023)).
@@
2.2 Models Structure and Types
@@
The type of those numbers is a key factor in the ability to compress the model parameters. Parameters typically represent real numbers and as such the most straightforward standard approach is to represent them with floating point numbers. The floating point format represents a flexible scale with a fixed number of bits which allows for larger numerical ranges. In a nutshell, floating point contains an exponent part - indicating the range in which the real number lies, a mantissa or fraction pointing to the actual number within this range, and a sign bit denoting whether a number is positive. For example, FP32 is a 32 bit floating point number with a sign bit, an 8 bit exponent and a 23 bit mantissa (see Figure 1). The real number is calculated by $(-1)^{sign}\cdot 2^{exponent-127}\cdot 1.fraction$. Many models are trained with FP32 for its high precision or a mix of precisions. Later during inference, the models sometimes keep their original format, but often choose to forgo some precision for a more compact representation and more efficient computation (Shoeybi et al., 2019; Granite Team, 2024; Dubey et al., 2024). Perhaps the most popular parameter type used for inference models is BF16 Wang & Kanwar (2019) which cuts the tail end of the fraction (hence reducing the precision level) but maintains the same exponent as shown in Figure 1. Other options include FP16, or quantized models using integer parameters with fixed precision.
2.3 Lossless compression
Lossless compressors are the traditional form of compression and are widely used for reducing network and storage overheads in all fields of computing. They consist of two algorithms – compression and decompression, where after applying both in sequence the output returns to the exact same state. There are countless compression techniques and they vary in the tradeoff between compressibility and compression/decompression time (see for example Squash (2016)).
@@ 2.3 Lossless compression
The main techniques employed in lossless compression are based on repetition removal (stemming from the seminal works of Lempel and Ziv Ziv & Lempel (1977; 1978) and dubbed LZ compression) and entropy encoding (e.g. Huffman (1952); Rissanen (1976)). LZ compressors find multiple-byte repetitions (typically of at least 4 bytes) and replace these with shorter back-pointers, hence saving space. Entropy encoding, on the other hand, looks at the entropy seen in the distribution of single bytes and reduces the length by working at a bit-level granularity, coding popular bytes with short representations. The most popular compressors combine the two techniques (first repetition removal and then entropy encoding), for example Zlib Adler & Gailly (2024) and Zstd Collet (2024b). [^2] In our experiments, we chose Zstd as the underlying compressor/decompressor due to its superior speed vs. compression tradeoff Collet (2024b); Squash (2016). A second family of compressors, that favors speed over compression ratio, relies solely on LZ repetition removal, e.g. LZ4 or Snappy Collet (2024a); Dean et al. (2024).
3 Compression for Models
In this section, we introduce ZipNN – our variants of lossless compression specifically tailored for model compression. We will first discuss compression of the majority of models and model contents. As mentioned in the introduction, we classify models into two categories - regular models and what we call clean models. Clean models are models that have undergone various techniques like number rounding or transformation between parameter types. These transformations typically leave many bits as zeros and increase model compressibility. However, once such a model is re-trained or fine tuned, it quickly becomes less compressible and behaves like a regular model. Regular models form the majority of models, and are models that were trained and remained unmodified after the training phase. In addition, we note that model formats also contain some metadata and at times a few layers that behave differently (for example tensors of parameters that contain integers). These typically constitute a negligible part of the model and hardly affect the model compression ratio. Later in the section, we will discuss clean models and variations.
3.1 Regular Model Compressibility
@@
A deeper dive shows that models indeed contain a high level of randomness in their parameters, but this is true for the fraction and sign bit parts of the floating point parameter. The exponent on the other hand is highly skewed and is the source of the compressibility. Figure 2 shows the distribution of the exponent in four different models. This distribution is highly skewed, and strikingly similar between different models, whether BF16 or FP32. It is very similar among language models but also close when testing other model types, such as Resnet which is an image model. Out of the 256 possible values in the exponent, we see only around 40 that actually appear (50 in the image model). Moreover, the top 12 values account for almost 99.9% of all parameters (17 in the image model).
@@
This highly skewed distribution can be explained as an artifact of the way models are trained. Model weights are initially set in the space of [-1,+1] and the training process scarcely pushes weights out of this range, perhaps since weights can have sufficient impact without the need to grow substantially. This explains why exponent values do not tend to exceed 128 which translates to the range [-1,1]. On the other hand, we see that weights are also not found in the lower ranges of extremely high precision. This is likely a result of various choices in the training process, for instance, optimizers (e.g. the Adam optimizer) add some noise in order to avoid division by zero. This noise is set in the order of $2^{-23}$ (corresponding to an exponent of around 99), and thus weights do not fall much below this noise level.
@@
In addition, we observe that there is no structure in the tensors, namely that the exponents are not ordered in any way. Whatever repetitions are found by an LZ type compression, these are most likely ”random” in the sense that they are an artifact of the skewed distribution. We therefore forgo the LZ compression part and use only an entropy encoder and specifically we use Huffman encoding. This proved to be helpful both in terms of compression/decompression performance, which is expected, but also in improving the compression ratio. It turns out that the random repetitions found by the LZ phase are naturally short, and not only do they not save much, but they also interfere with the effectiveness of the Huffman encoder (by adding back-pointers to the mix). In order to ensure that the repetitions found are indeed ”random”, we ran a test by which we do a random shuffle of the parameters in a model and then compress the exponents using Zstd. The shuffled version reached nearly the same compression ratio (up to 0.05%). Our experiments show that using Huffman encoding gives a very good trade-off of time vs. space. An FSE entropy encoder achieves a slightly better compression ratio (0-2%) at the expense of significant performance penalty (at times over 2X).
Figure 4 shows the compression ratio benefits of each of the steps that we use. We see that using Huffman without exponent extraction is only helpful for speed over Zstd. However, once we divide the exponent from the rest of the data, using Huffman vs Zstd also achieves better compression (as well as speed).
@@
When compressing regular models we found that the exponent compresses by approximately a 3X factor (namely compressed size of 33%) whereas the fraction and sign bits hardly compress at all. The overall savings is thus dictated by the weight of the exponent in the parameter. That is, for BF16 the compressed size is $\frac{1}{2}\cdot 33.3+\frac{1}{2}\cdot 100\approx 66.6$, whereas for FP32 it is $\frac{1}{4}\cdot 33.3+\frac{3}{4}\cdot 100\approx 83.3$. As a result, the main benefit of compressing regular models lies in BF16 models (which fortunately are most of the new models used in practice). We show the effect on performance of the exponent extraction and Huffman only encoding in Section 5.2
3.2 Compressing Clean Models
As mentioned earlier, most models only show compressibility in the exponent. However, we also see some models that are surprisingly more compressible, and show compressibility also in the fraction part of the parameters. We call these clean models, since after further fine tuning they typically lose their extra compressibility. That being said, these include some very popular models, such as Bge, RoBERTa, Xlm-RoBERTa, Whisper, and Clip.

### v2

@@ Abstract
We present ZipNN, a lossless compression tailored to neural networks. Somewhat surprisingly, we show that specific lossless compression can gain significant network and storage reduction on popular models, often saving $33\%$ and at times reducing over $50\%$ of the model size. We investigate the source of model compressibility and introduce specialized compression variants tailored for models that further increase the effectiveness of compression. On popular models (e.g. Llama 3) ZipNN shows space savings that are over $17\%$ better than vanilla compression while also improving compression and decompression speeds by $62\%$. Using multiple workers and threads, ZipNN can achieve decompression speeds of up to 80GB/s and compression speed of up to 13GB/s. We estimate that these methods could save over an ExaByte per year of network traffic downloaded from a large model hub like Hugging Face.
Index Terms:
compression, lossless compression, models, AI, language models
I Introduction
With scale, we have learned that models gain stronger abilities and with it, gain popularity. With scale, models also require more storage space and memory, and with popularity, more communication bandwidth. Taken together, we observe strains on communication bottlenecks that call for efficient solutions. Storage requirements, while often ignored, may accumulate to hundreds or thousands of times the size of a model if checkpoints [1] or distributed updates are to be saved (c.f., VI) [2, 3, 4].
Similarly, models are repeatedly moved around in multiple channels: from a storage hub to inference machines; from training/fine-tuning nodes to the storage backend; between GPU nodes during distributed training, and so on. Network hubs epitomize the strains by model size. For instance, with over 14.5 GBs and 2.77 M downloads per month from Hugging Face [5] Mistral [6] alone requires 40 PBs of transferred information a month.
A large body of work has been aimed at reducing model sizes focusing on the parameter sizes and the number of computations in inference. Such methods transform the model into a smaller one in an irreversible fashion. For example, distillation [7], pruning [8] and quantization [9] all lose some information and, potentially, performance, for improved efficiency. Since these methods’ main focus is on inference speed, they are bound to the format of an actual running model. As such, they don’t necessarily push the space-saving to its limits, and are not stored in the minimal possible way.
@@ I Introduction
Surprisingly, we observe (§III) that even standard lossless compressors like zlib [10] or zstd [11] can achieve non-negligible savings. While common rationale expects model parameters to have high entropy and therefore be non-compressible, we find that in reality there is less entropy than what the representation offers. Our goal is therefore to find a compression method that maximizes compression benefits for models while also being aware of compression and decompression speeds.
We identify the source of model compressibility as the floating point range that actually exists in models. Specifically, we find that the exponent component in a floating point parameter is highly skewed and therefore highly compressible. To this end, we devise a compression method that separates the exponent bits from the rest of the bits that often show no compressibility. We also identified that the source of compressibility of the exponent is entirely the skewed distribution of single bytes. This means that compressors that search for multi-byte repetitions are both unnecessarily time-consuming and unhelpful in this case, and in particular all of the Lempel-Ziv algorithms [12, 13] hardly achieve any data reduction. Instead, we use only Entropy Encoding and specifically Huffman codes [14], improving both performance and compression ratio in doing so. Another speed improvement comes from identifying the non-compressible parts and avoiding time-consuming compression attempts on them. The observations described above are relevant to most of the models, but not all of them. ZipNN therefore identifies the characteristics of the model at hand and determines which strategy best suits it.
@@
We classify popular models into categories with distinct compressibility traits. We highlight the clean models category, where models have undergone some rounding after the training phase. The rounding allows compression of the other bits (apart from the exponent) and therefore clean models achieve effective compression, at times reaching 55% savings. For Example, the popular RoBERTa model [15] falls into this category. Our compressor is therefore challenged with the task of inline identifying the potential space savings in such a model and compressing it accordingly.
When comparing ZipNN to a state-of-the art compressor like Zstd, we get the following improvements: On BF16 models our compressor achieves a 17% improvement in compression ratio and a 62% speedup in compression/decompression. For clean models, this is even larger with a 34% improvement in space and a 4.6X speedup in compression time and 83% speedup in decompression. By using multi-threading we show that ZipNN can reach a throughput of up to 80GB/s for decompression and up to 13GB/s for compression.
@@
II Background
II-A Motivation - use cases
With small models weighing about a Gigabyte [16] and large ones Terrabytes [17] storage and network overheads become an issue for many purposes. Moreover, common use cases require many model types or model versions and hence increased resources. We list some below as a motivation.
TABLE I: Top ranked downloaded models from Hugging Face.
@@ II-A1 Model Hubs
Large model repositories or hubs like Hugging Face [5], Model Zoo [18], PyTorch [19], Tensorflow [20], Adapter [21], IBM watsonx.data [22] and Qualcomm® AI [23] hold a large number of models and serve numerous download requests of popular models. Hugging Face, the largest of these hubs, revealed in a statement in August 2024, that it holds 1.3M models, with a cumulative storage space of 12PB. They also serve 1 billion daily requests amounting to a network bandwidth of around 6 PetaBytes per day! Table I shows some of the top-ranked models [^1] Rank as of October 2024. and their compression ratio (using the methods described in Section II-C). As seen, the potential traffic savings from compression is substantial.
@@
In this use case, there are three ways in which compression can be beneficial, the first is to reduce the amount of data transferred, the second is to reduce the time to download and upload models and the third is to reduce the amount of data stored. In a cloud based hub like Hugging Face, which runs over the internet, the network benefits are crucial (in particular the first one) (see Section V-D). On the other hand, in use-cases of serving models with high performance, pricey storage, typically the storage savings becomes more important (e.g. see Section V-D2).
II-A2 Distributed and Decentralized Training
@@
During training of large models, training nodes transfer data between them to overcome the need to save the full model and computation on a single GPU/node. In some methods, only the model weights are transferred between nodes, and in other methods, the optimizer weights and gradients are transferred as well [24]. Either way, distributed training is usually limited by data transfer between nodes, an issue compression addresses directly. For example: FSDP (Fully Sharded Data Parallel) is a PyTorch method that during training, in addition to transferring model weights, also transfers gradients and optimizer states, causing a networking bottleneck that limits the training of large models.
@@
Another training paradigm is decentralized or federated training, which proposes accommodating training of the same model by different contributors. This ranges from federated learning that contributes gradients [4], to contributing partially trained models [25], from changing the kinds of updates done [26] to relying on volunteer computing [27], or even relying on different objectives and expertise all merged into the same model [3]. All of these methods inherently transfer and store numerous model versions, which also drove dedicated version control frameworks [2].
II-A3 Checkpoints and Versions
@@
During model creation, multiple intermediate versions of the models are commonly saved. This often includes tests on the training regime such as hyperparameter tuning [28]. Even during the training of a single model, the current model is periodically checkpointed to recover after a crash, to select the best checkpoint from a few options [29], for analysis [1], improve performance [30, 31], etc. Even though saving during checkpointing rarely slows the training time, it does burden the storage (as well as the network), limiting the frequency and amount of saved and shared checkpoints, which are encouraged by the community (e.g.; [1, 32]).
@@
II-B Models Structure and Types
@@
The type of those numbers is a key factor in the ability to compress the model parameters. Parameters typically represent real numbers and as such the most straightforward standard approach is to represent them with floating point numbers. The floating point format represents a flexible scale with a fixed number of bits which allows for larger numerical ranges. In a nutshell, floating point contains an exponent part - indicating the range in which the real number lies, a mantissa or fraction pointing to the actual number within this range, and a sign bit denoting whether a number is positive. For example, FP32 is a 32 bit floating point number with a sign bit, an 8 bit exponent and a 23 bit mantissa (see Figure 1). The real number is calculated by $(-1)^{sign}\cdot 2^{exponent-127}\cdot 1.fraction$. Many models are trained with FP32 for its high precision or a mix of precisions. Later during inference, the models sometimes keep their original format, but often choose to forgo some precision for a more compact representation and more efficient computation [33, 34, 35]. Perhaps the most popular parameter type used for inference models is BF16 [36] which cuts the tail end of the fraction (hence reducing the precision level) but maintains the same exponent as shown in Figure 1. Other options include FP16, or quantized models using integer parameters with fixed precision.
II-C Lossless compression
Lossless compressors are the traditional form of compression and are widely used for reducing network and storage overheads in all fields of computing. They consist of two algorithms – compression and decompression, where after applying both in sequence the output returns to the exact same state. There are countless compression techniques and they vary in the tradeoff between compressibility and compression/decompression time (see for example [37]).
@@ II-C Lossless compression
The main techniques employed in lossless compression are based on repetition removal (stemming from the seminal works of Lempel and Ziv [12, 13] and dubbed LZ compression) and entropy encoding (e.g. [14, 38]). LZ compressors find multiple-byte repetitions (typically of at least 4 bytes) and replace these with shorter back-pointers, hence saving space. Entropy encoding, on the other hand, looks at the entropy seen in the distribution of single bytes and reduces the length by working at a bit-level granularity, coding popular bytes with short representations. The most popular compressors combine the two techniques (first repetition removal and then entropy encoding), for example Zlib [39] and Zstd [40]. [^2] In our experiments, we chose Zstd as the underlying compressor/decompressor due to its superior speed vs. compression tradeoff [40, 37]. A second family of compressors, that favors speed over compression ratio, relies solely on LZ repetition removal, e.g. LZ4 or Snappy [41, 42].
III Compression for Models
In this section, we introduce ZipNN – lossless compression specifically tailored for model compression. We will first discuss compression of the majority of models and model contents. As mentioned in the introduction, we classify models into two categories - regular models and what we call clean models. Clean models are models that have undergone various techniques like number rounding or transformation between parameter types. These transformations typically leave many bits as zeros and increase model compressibility. However, once such a model is re-trained or fine tuned, it quickly becomes less compressible and behaves like a regular model. Regular models form the majority of models, and are models that were trained and remained unmodified after the training phase. In addition, we note that model formats also contain some metadata and at times a few layers that behave differently (for example tensors of parameters that contain integers). These typically constitute a negligible part of the model and hardly affect the model compression ratio. Later in the section, we will discuss clean models and variations.
III-A Regular Model Compressibility
@@
A deeper dive shows that models indeed contain a high level of randomness in their parameters, but this is true for the fraction and sign bit parts of the floating point parameter. The exponent on the other hand is highly skewed and is the source of the compressibility. Figure 2 shows the distribution of the exponent in four different models. This distribution is highly skewed, and strikingly similar between different models, whether BF16 or FP32. It is very similar among language models, but also close when testing other model types, such as Resnet, which is an image model. Out of the 256 possible values in the exponent, we see only around 40 that actually appear (50 in the image model). Moreover, the top 12 values account for almost 99.9% of all parameters (17 in the image model).
@@
This highly skewed distribution can be explained as an artifact of the way models are trained. Model weights are initially set in the space of [-1,+1] and the training process scarcely pushes weights out of this range, perhaps since weights can have sufficient impact without the need to grow substantially. This explains why exponent values do not tend to exceed 128, which translates to the range [-1,1]. On the other hand, we see that weights are also not found in the lower ranges of extremely high precision. This is likely a result of various choices in the training process, for instance, optimizers (e.g. the Adam optimizer) add some noise in order to avoid division by zero. This noise is set in the order of $2^{-23}$ (corresponding to an exponent of around 99), and thus weights do not fall much below this noise level.
@@
In addition, we observe that there is no structure in the tensors, namely that the exponents are not ordered in any way. Whatever repetitions are found by an LZ type compression, these are most likely ”random” in the sense that they are an artifact of the skewed distribution. We therefore forgo the LZ compression part and use only an entropy encoder and specifically we use Huffman encoding. This proved to be helpful both in terms of compression/decompression performance, which is expected, but also in improving the compression ratio. It turns out that the random repetitions found by the LZ phase are naturally short, and not only do they not save much, but they also interfere with the effectiveness of the Huffman encoder (by adding back-pointers to the mix). In order to ensure that the repetitions found are indeed ”random”, we shuffled the parameters in a model randomly and then compressed the exponents using Zstd. The shuffled version reached nearly the same compression ratio (up to 0.05%). This experiment shows that Huffman encoding provides a very good trade-off of time vs. space. An FSE entropy encoder achieves a slightly better compression ratio (0-2%) at the expense of significant performance penalty (at times over 2X).
Figure 4 shows the compression ratio benefits of each of our steps. We see that using Huffman without exponent extraction is only helpful for speed over Zstd. However, once we divide the exponent from the rest of the data, using Huffman vs Zstd also achieves better compression (as well as speed).
@@
When compressing regular models we found that the exponent compresses by approximately a 3X factor (namely compressed size of 33%) whereas the fraction and sign bits hardly compress at all. The overall savings is thus dictated by the weight of the exponent in the parameter. That is, for BF16 the compressed size is $\frac{1}{2}\cdot 33.3+\frac{1}{2}\cdot 100\approx 66.6$, whereas for FP32 it is $\frac{1}{4}\cdot 33.3+\frac{3}{4}\cdot 100\approx 83.3$. As a result, the main benefit of compressing regular models lies in BF16 models (which fortunately are most of the new models used in practice). We show the effect on performance of the exponent extraction and Huffman only encoding in Section V-B
III-B Compressing Clean Models
As mentioned earlier, most models only show compressibility in the exponent. However, we also see some models that are surprisingly more compressible, and show compressibility also in the fraction part of the parameters. We call these clean models, since after further fine tuning they typically lose their extra compressibility. That being said, these include some of the most popular models, such as Bge, RoBERTa, Xlm-RoBERTa, Whisper, and Clip.